Find the co-efficient of correlation between persons employed(X) and cloth manufactured(Y) in a textile Mill. Interpret the results.                                                     

X

137

209

113

189

176

200

219

Y

23

47

22

40

39

51

49


1
Expert's answer
2021-07-07T17:31:27-0400

"r =
 \frac{ \sum_{i=1}^{n}(x_i-\bar{x})(y_i-\bar{y}) }{%
 \sqrt{\sum_{i=1}^{n}(x_i-\bar{x})^2}\sqrt{\sum_{i=1}^{n}(y_i-\bar{y})^2}}"





From the table above;

"r =
 \frac{ 2693.142857 }{
 \sqrt{9155.71429}\times \sqrt {853.428571}}"

"=0.96345"

The coefficient of correlation between persons employed(X) and cloth manufactured(Y) in a textile Mill is 0.96345. There is a strong positive association between persons employed and cloth manufactured.